OpenStudy (anonymous):

solve for x..... 2^((x^2)-x)=64

OpenStudy (anonymous):

x = 3 2 ^ ((3^2)-3) = 64 2^(9 -3) = 64 2^6 = 64 2*2*2*2*2*2 = 64 64 = 64

OpenStudy (anonymous):

but how did you find 3...or did you just plug in random numbers??

OpenStudy (anonymous):

Usually in these types of homework exercises the answer is a small positive integer. So I just plugged in some numbers, you can use logarithms if you find that better. log base 2 of 2 ^((x^2) -x) = log base 2 of 64 x^2-x=6 x(x-1)=6 x=3 x= -2

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

can you explain?

OpenStudy (anonymous):

You apply log base 2 to both sides. Do you know how to use log? log base x of y = b means that x^b = y log base 10 of 100 = 2 because 10^2 = 100.

OpenStudy (anonymous):

yes i know that...

OpenStudy (anonymous):

nevermind i understand...thankyou!

OpenStudy (anonymous):

You're welcome.